# Class-Dependent Perturbation Effects in Evaluating Time Series Attributions

### v1

@@ Abstract
As machine learning models become increasingly prevalent in time series applications, Explainable Artificial Intelligence (XAI) methods are essential for understanding their predictions. Within XAI, feature attribution methods aim to identify which input features contributed the most to a model’s prediction, with their evaluation typically relying on perturbation-based metrics. Through empirical analysis across multiple datasets, model architectures, and perturbation strategies, we identify important class-dependent effects in these metrics: they show varying effectiveness across classes, achieving strong results for some while remaining less sensitive to others. In particular, we find that the most effective perturbation strategies often demonstrate the most pronounced class differences. Our analysis suggests that these effects arise from the learned biases of classifiers, indicating that perturbation-based evaluation may reflect specific model behaviors rather than intrinsic attribution quality. We propose an evaluation framework with a class-aware penalty term to help assess and account for these effects in evaluating feature attributions. Although our analysis focuses on time series classification, these class-dependent effects likely extend to other structured data domains where perturbation-based evaluation is common. [^1] Code and results are available at https://github.com/gregorbaer/class-perturbation-effects. .
@@ 2 Related Work
We focus on perturbation analysis as a computational method to measure the correctness and sparsity of explanations. This approach was first introduced by Samek et al. [12] to evaluate feature attribution methods in the image domain. It involves sequentially perturbing pixels in order of the most relevant features first by replacing them with noninformative values and observing the impact on model predictions. This process generates a perturbation curve that tracks these prediction changes, allowing the calculation of metrics such as the area under or over the curve to jointly measure the correctness and sparsity of explanations. The fundamental assumption underlying this approach is that perturbing important features should degrade model predictions proportionally to their attributed importance, while perturbing irrelevant features should have minimal effects on the model output.
@@
Schlegel et al. [13] were the first to apply perturbation analysis to time series classification, evaluating the quality of attributions by measuring the average change in accuracy over the perturbed samples with four perturbation strategies, including zero and mean value replacement. Mercier et al. [8] expanded on this framework by evaluating attributions with additional metrics from image explanations, such as sensitivity and infidelity, revealing that no single attribution method consistently outperforms others in all aspects of evaluation.
@@
However, an important question has remained unexplored: how do the characteristics of different classes affect perturbation-based evaluation methods? Perturbation strategies may be influenced by classifier biases. For example, if a classifier learned to associate certain perturbation values (such as zero) with specific classes, the effectiveness of perturbation-based evaluation could vary between different predicted classes. This consideration may help explain some findings in the literature. For example, Schlegel et al. [14] evaluated 16 different perturbation approaches and found that for most datasets, only up to 60% of the samples changed their predicted label under perturbation, regardless of the strategy employed. Such results suggest that the effectiveness of perturbations might be influenced by factors beyond the perturbation strategy itself. Our work investigates whether class-dependent effects could explain these observed patterns, examining how the relationship between perturbation strategies and learned class representations might affect evaluation outcomes.
@@ 3 Class-adjusted Perturbation Analysis
Although we apply this penalty framework to the mean DS, it generalizes to any evaluation metric. This allows a systematic investigation of class-dependent effects in attribution evaluation by comparing base metrics against their class-adjusted variants. We use this framework to analyze how perturbation-based evaluation methods may exhibit different behaviors across classes.
@@ 4 Experiment Set-Up
We use four univariate time series datasets from the UCR Time Series Classification Archive [1]: FordA, FordB, ElectricDevices (ElecDev) and Wafer. Table 1 presents the characteristics of these datasets along with the achieved accuracy of the classifiers used in this study. These datasets are among the largest in the UCR Archive and are commonly used to evaluate feature attributions with perturbations [13, 18, 8, 14, 15, 23, 8], making our results comparable to previous work.
@@
We evaluate five widely adopted attribution methods. [^2] We implement all attribution methods using the TSInterpret package [5]. These include four gradient-based methods: Gradients (GR) [19], Integrated Gradients (IG) [21], SmoothGrad (SG) [20], and Gradient SHAP (GS) [7]. We also include a method based on perturbation, Feature Occlusion (FO) [4]. To ensure balanced class representation while maintaining computational feasibility, we compute attributions on a stratified sample of 300 instances per class from the test set. Attributions always explain the predicted class label.
@@ 5.1 Evaluating Attribution Quality
Table 3: Mean degradation scores ($\overline{\text{DS}}$) for named perturbation strategies across datasets, models, and attribution methods. $\overline{\text{DS}}$ measures the average differential impact between perturbing most and least relevant features. Positive values indicate correct feature identification, values near zero suggest non-discriminative attributions, and negative values indicate reversed feature importance. Bold and italic values highlight the highest and second-highest values per dataset and model-attribution combination.
@@
FordA Gauss -0.01 0.06 0.02 -0.04 0.09 -0.00 0.11 0.01 -0.02 0.14
Inv -0.00 0.01 0.01 0.01 0.02 -0.00 0.02 0.00 0.01 0.02
Opp -0.03 0.03 0.01 0.03 0.04 -0.05 0.06 0.02 0.04 0.06
SubMean -0.04 0.07 0.03 0.07 0.13 -0.09 0.17 0.05 0.14 0.20
Unif -0.01 0.01 0.00 0.01 0.01 -0.01 0.02 0.01 0.02 0.02
Zero -0.04 0.07 0.03 0.07 0.13 -0.10 0.17 0.05 0.15 0.20
FordB Gauss -0.00 0.02 -0.00 -0.03 0.03 -0.03 0.13 0.01 -0.03 0.18
Inv -0.00 0.00 -0.00 0.00 0.00 -0.01 0.02 0.01 0.01 0.02
Opp -0.01 0.02 0.00 0.02 0.02 -0.04 0.03 -0.02 0.03 0.03
SubMean -0.02 0.04 0.01 0.03 0.06 -0.07 0.09 -0.07 0.08 0.10
Unif -0.00 0.00 0.00 0.00 0.01 -0.01 0.01 0.00 0.00 0.01
Zero -0.02 0.04 0.01 0.03 0.06 -0.07 0.09 -0.07 0.08 0.10
Wafer Gauss -0.06 -0.10 0.01 0.08 -0.16 0.02 -0.10 0.12 0.18 -0.20
Inv 0.02 0.06 0.04 0.04 0.14 -0.01 0.09 0.03 0.07 0.24
Opp -0.02 0.01 0.01 0.01 0.02 -0.01 0.06 0.04 0.05 0.13
SubMean 0.12 0.20 0.01 0.17 0.33 0.06 0.05 0.03 0.07 0.06
Unif -0.01 0.01 0.03 0.02 0.03 0.03 0.02 0.07 0.02 0.10
Zero -0.02 0.06 0.05 0.05 0.14 0.02 0.08 0.04 0.05 0.28
ElecDev Gauss -0.00 -0.01 0.01 0.01 -0.01 0.03 -0.00 0.08 0.05 -0.04
Inv 0.01 0.06 0.08 0.05 0.07 -0.03 0.06 0.01 0.06 0.03
Opp 0.02 0.18 0.06 0.14 0.22 0.07 0.18 0.09 0.18 0.20
SubMean 0.14 0.24 0.22 0.22 0.27 0.15 0.22 0.11 0.20 0.29
Unif 0.05 0.09 0.05 0.06 0.16 0.04 0.10 0.05 0.07 0.11
Zero 0.04 0.16 0.08 0.12 0.27 0.09 0.19 0.08 0.16 0.27
Examining these results in detail, we find that the perturbation strategies exhibit varying levels of performance. Zero and SubMean perturbations consistently achieve superior results across most experimental configurations, ranking among the top performers in terms of mean DS. This suggests that they are generally effective across different datasets and model architectures. However, this pattern has notable exceptions. For the InceptionTime classifier, Gauss perturbations demonstrate superior performance in two specific cases. On the FordB dataset, they achieve the highest performance with IG, SG, FO attributions ($\overline{\text{DS}}$ equal to 0.13, 0.01 and 0.18, respectively). Similarly, on the Wafer dataset, they excel with SG and GS attributions ($\overline{\text{DS}}$ equal to 0.12 and 0.18, respectively). In all other configurations, Gauss perturbations remain less effective.
Turning to attribution methods, we observe substantial performance variations. FO consistently achieves the highest DS values across datasets and perturbation strategies, indicating superior feature importance identification. IG and GS demonstrate comparable performance levels, though generally lower than FO. In contrast, GR shows poor discriminative ability, frequently yielding negative or near-zero DS values, suggesting that its feature importance assignments are often not better than random ordering.
@@ 5.2 Distribution Patterns in Attribution Quality
(a) Overall DS distributions per attribution method.
(b) Class-specific DS distributions per attribution method.
@@
To understand how attribution performance varies between individual instances, we analyze the distributional characteristics of DS scores. We select the FordB dataset, InceptionTime architecture, and SubMean perturbation strategy as a case study, as these demonstrate patterns typical of our broader findings. Figure 3 presents these distributions, revealing both overall performance patterns and, crucially, class-specific effects.
The aggregate distributions shown in Figure 3(a) reveal notable variation in attribution method effectiveness. IG, GS, and FO exhibit positively skewed distributions with extended tails toward higher DS values, indicating superior attribution quality. In contrast, GR and SG demonstrate negative skewness with tails extending toward lower DS values, suggesting less reliable feature identification. A critical observation is the concentration of scores around zero across all methods, with mean DS values confined to a narrow range of [-0.07, 0.10] also observed in Table 3. This clustering implies limited discriminative power in feature ordering for many instances, even among the better-performing methods.
Class-stratified analysis, presented in Figure 3(b), uncovers substantial heterogeneity in attribution quality across classes. For IG, GS, and FO, instances from class 1 consistently achieve higher DS scores with pronounced positive tails and minimal negative values. However, these same methods show markedly different behavior for class 0, where DS scores cluster tightly around zero, indicating minimal perturbation impact. GR and SG show an inverse pattern, with class 1 showing predominantly negative scores while class 0 maintains a more balanced distribution centered near zero.
These pronounced class-dependent variations in attribution performance raise questions about the generalizability of perturbation-based evaluation methods. To establish whether these observed patterns represent a systematic phenomenon beyond this specific case study, we extend our investigation to examine class-dependent behavior across all experimental conditions.
@@ 5.3 Class-Dependent Effects in Perturbation Analysis
To systematically investigate the presence and extent of class-dependent effects in perturbation-based evaluation, we apply the proposed class-adjusted DS ($\text{DS}_{c}$) metric to all experimental conditions. Table 4 presents these adjusted scores, which balance average attribution correctness with consistency between classes. A notable pattern emerges: the incorporation of class consistency penalties substantially reduces performance metrics across most experimental conditions, particularly for previously high-performing perturbation strategies.
Table 4: Class-adjusted mean degradation scores ($\text{DS}_{c}$ with $\alpha=1$) for named perturbation strategies across datasets, models, and attribution methods. $\text{DS}_{c}$ measures both attribution quality and cross-class consistency of perturbations. Positive values indicate good attribution quality with consistent behavior across classes, while lower values suggest either poor attribution quality or inconsistent performance between classes. Bold and italic values highlight the highest and second-highest values per dataset and model-attribution combination.
@@
FordA Gauss -0.01 0.00 0.00 -0.08 0.00 -0.01 0.00 -0.00 -0.05 0.00
Inv -0.00 0.00 0.00 0.00 0.00 -0.01 0.00 -0.00 0.00 0.00
Opp -0.05 0.00 0.00 0.00 0.00 -0.10 0.00 -0.00 0.00 0.00
SubMean -0.08 0.00 0.00 0.00 0.01 -0.19 0.00 0.00 0.00 -0.00
Unif -0.02 0.00 0.00 0.00 0.00 -0.03 -0.00 -0.00 0.00 -0.00
Zero -0.08 0.00 0.00 0.00 0.01 -0.20 0.00 0.00 -0.00 -0.00
FordB Gauss -0.00 0.00 -0.00 -0.05 0.01 -0.07 0.01 0.01 -0.05 0.03
Inv -0.00 0.00 -0.00 0.00 0.00 -0.01 0.00 0.00 0.00 0.00
Opp -0.02 0.00 0.00 0.00 0.00 -0.09 0.00 -0.04 0.00 0.01
SubMean -0.04 0.01 0.00 0.00 0.01 -0.16 0.01 -0.13 0.01 0.01
Unif -0.01 0.00 0.00 0.00 0.00 -0.02 0.00 0.00 0.00 0.00
Zero -0.04 0.01 0.00 0.01 0.01 -0.16 0.01 -0.13 0.01 0.02
Wafer Gauss -0.15 -0.21 0.00 0.00 -0.37 0.00 -0.20 -0.00 0.03 -0.42
Inv 0.01 0.00 -0.00 0.00 0.00 -0.03 0.00 -0.01 0.01 0.01
Opp -0.04 0.00 -0.00 0.00 0.00 -0.02 0.00 -0.01 0.01 0.01
SubMean 0.03 0.02 -0.01 0.04 0.04 0.01 0.05 -0.00 0.02 0.03
Unif -0.02 0.00 0.00 0.00 0.00 0.00 0.00 -0.01 0.01 0.01
Zero -0.05 0.01 -0.00 0.01 0.01 0.01 0.01 -0.02 0.04 0.02
ElecDev Gauss -0.04 -0.04 -0.02 -0.03 -0.04 0.01 -0.01 0.04 0.03 -0.09
Inv -0.02 0.01 0.05 0.01 0.03 -0.08 0.02 -0.05 0.03 -0.04
Opp -0.03 0.10 0.00 0.10 0.16 0.04 0.14 0.04 0.12 0.12
SubMean 0.08 0.13 0.14 0.14 0.22 0.07 0.15 0.08 0.13 0.23
Unif 0.03 0.03 -0.00 0.02 0.08 -0.00 0.06 0.03 0.04 0.02
Zero -0.01 0.09 0.03 0.06 0.23 0.05 0.12 0.04 0.10 0.16
@@
Zero and SubMean perturbation strategies, which demonstrated superior performance before, show marked degradation under class-adjusted evaluation. For the FordA and FordB datasets, these strategies’ scores approach zero or become negative after applying class-adjusted penalties, suggesting that their apparent effectiveness stems primarily from class-specific behavior rather than genuine attribution quality over all instances. The impact varies by dataset. Wafer and ElecDev show more resilience to class adjustment, though we still see reduced scores. For ElecDev, the interpretation of these results requires additional context: as the only multiclass dataset in our evaluation, the pairwise averaging of class differences may underestimate class-specific effects due to the higher dimensionality of the classification space.
@@
To examine the mechanisms underlying these class-specific effects, Figure 4 presents a detailed analysis of perturbation impacts across classes on the FordB dataset. Figure 4(a), which focuses on named perturbation strategies with InceptionTime and FO attribution, reveals systematic class-dependent behavior. Class 0 instances consistently demonstrate minimal response to perturbation across all strategies, with DS scores tightly clustered around zero. In contrast, Class 1 instances show substantial variability in perturbation response, with Gauss achieving highest effectiveness, followed by Zero and SubMean strategies. This asymmetric response pattern explains the earlier observed degradation in class-adjusted metrics: perturbation strategies succeed primarily by exploiting class-specific model behaviors.
@@
Analysing the constant-value perturbations in Figure 4(b) reinforces these findings while providing additional information. The evaluation of perturbation values between -2 and 2 reveals optimal effectiveness at moderate positive values, particularly around 0.5, challenging the conventional preference for zero-based perturbation. We now extend the investigation of the constant perturbation strategies to all experimental conditions.
(a) DS for named perturbation strategies.
(b) DS for constant perturbation strategies.

### v2

@@ Abstract
As machine learning models become increasingly prevalent in time series applications, Explainable Artificial Intelligence (XAI) methods are essential for understanding their predictions. Within XAI, feature attribution methods aim to identify which input features contribute the most to a model’s prediction, with their evaluation typically relying on perturbation-based metrics. Through systematic empirical analysis across multiple datasets, model architectures, and perturbation strategies, we reveal previously overlooked class-dependent effects in these metrics: they show varying effectiveness across classes, achieving strong results for some while remaining less sensitive to others. In particular, we find that the most effective perturbation strategies often demonstrate the most pronounced class differences. Our analysis suggests that these effects arise from the learned biases of classifiers, indicating that perturbation-based evaluation may reflect specific model behaviors rather than intrinsic attribution quality. We propose an evaluation framework with a class-aware penalty term to help assess and account for these effects in evaluating feature attributions, offering particular value for class-imbalanced datasets. Although our analysis focuses on time series classification, these class-dependent effects likely extend to other structured data domains where perturbation-based evaluation is common. [^1] Code and results are available at https://github.com/gregorbaer/class-perturbation-effects. .
@@ 2 Related Work
We focus on perturbation analysis as a computational method to measure the correctness and compactness of explanations, properties identified as key quality criteria for XAI methods [9]. This approach was first introduced by Samek et al. [12] to evaluate feature attribution methods in the image domain. It involves sequentially perturbing pixels in order of the most relevant features first by replacing them with noninformative values and observing the impact on model predictions. This process generates a perturbation curve that tracks these prediction changes, allowing the calculation of metrics such as the area under or over the curve to jointly measure the correctness and compactness of explanations.
In this context, correctness refers to whether an explanation faithfully identifies features that truly influence the model’s prediction. Compactness captures how concisely the explanation represents the model’s behavior. A compact explanation would show a quick degradation of predictions when perturbing just a few highly-ranked features, indicating that the model relies on a small subset of input features. The fundamental assumption underlying this approach is that perturbing important features should degrade model predictions proportionally to their attributed importance, while perturbing irrelevant features should have minimal effects on the model output.
@@
Figure 2 demonstrates how feature attributions can guide the perturbation process, where important time points or segments of a time series (indicated by darker red) are replaced with non-informative values like zero. This approach forms the basis of perturbation-based evaluation methods discussed below. Schlegel et al. [13] were the first to apply perturbation analysis to time series classification, evaluating the quality of attributions by measuring the average change in accuracy over the perturbed samples with four perturbation strategies, including zero and mean value replacement. Mercier et al. [8] expanded on this framework by evaluating attributions with additional metrics from image explanations, such as sensitivity and infidelity, revealing that no single attribution method consistently outperforms others in all aspects of evaluation.
Figure 2: Schematic illustration of time series perturbation based on feature attributions. Reddish rectangles represent feature attribution importance (darker red indicates higher importance). This example shows one perturbation approach: replacing values in the most important region with a constant value (e.g. zero).
@@
However, an important question has remained unexplored: how do the characteristics of different classes affect perturbation-based evaluation methods? Perturbation strategies may be influenced by classifier biases. For example, if a classifier learned to associate certain perturbation values (such as zero) with specific classes, the effectiveness of perturbation-based evaluation could vary between different predicted classes. This phenomenon occurs when perturbation values inadvertently match features the model has associated with a specific class. As a result, substituting “important” features with these values might paradoxically reinforce rather than disrupt the prediction, regardless of attribution correctness.
This consideration may help explain some findings in the literature. For example, Schlegel et al. [14] evaluated 16 different perturbation approaches and found that for most datasets, only up to 60% of the samples changed their predicted label under perturbation, regardless of the strategy employed. These results suggest that the effectiveness of perturbations might be influenced by factors beyond the perturbation strategy itself. Our work investigates whether class-dependent effects could explain these observed patterns, examining how the relationship between perturbation strategies and learned class representations might affect evaluation outcomes.
@@ 3 Class-Adjusted Perturbation Analysis
Although we apply this penalty framework to the mean DS, it generalizes to any evaluation metric. This allows a systematic investigation of class-dependent effects in attribution evaluation by comparing base metrics against their class-adjusted variants while requiring minimal additional computation, as it utilizes class information already collected during the standard perturbation process. We use this framework to analyze how perturbation-based evaluation methods can exhibit different behaviors across classes.
@@ 4 Experiment Set-up
We use four univariate time series datasets from the UCR Time Series Classification Archive [1]: FordA, FordB, ElectricDevices (ElecDev) and Wafer. Table 1 presents the characteristics of these datasets along with the achieved accuracy of the classifiers used in this study. For visualizations of representative samples and detailed class distributions, see Appendix 0.A. These datasets are among the largest in the UCR Archive and are commonly used to evaluate feature attributions with perturbations [13, 18, 8, 14, 15, 23, 8], making our results comparable to previous work.
@@
We evaluate five widely adopted attribution methods. [^2] We implement all attribution methods using the TSInterpret package [5]. These include four gradient-based methods: Gradients (GR) [19], Integrated Gradients (IG) [21], SmoothGrad (SG) [20], and Gradient SHAP (GS) [7]. We also include a method based on perturbation, Feature Occlusion (FO) [4]. While gradient-based methods compute feature importance by analyzing how changes in inputs affect model outputs through gradient calculations, perturbation-based methods systematically modify input features and observe the resulting changes in predictions. To ensure balanced class representation while maintaining computational feasibility, we compute attributions on a stratified sample of 300 instances per class from the test set. Attributions always explain the predicted class label.
@@ 5.1 Evaluating Attribution Quality
Table 3: Mean degradation scores ($\overline{\text{DS}}$) for named perturbation strategies across datasets, models, and attribution methods. $\overline{\text{DS}}$ measures the average differential impact between perturbing most and least relevant features. Positive values indicate correct feature identification, values near zero suggest non-discriminative attributions, and negative values indicate reversed feature importance. The largest values for each dataset and model-attribution column pair are highlighted in bold.
@@
FordA Gauss -.006 .058 .021 -.040 .088 -.004 .109 .012 -.025 .141
Inv -.002 .014 .008 .012 .015 -.004 .016 .001 .011 .018
Opp -.027 .031 .011 .027 .040 -.052 .058 .018 .044 .061
SubMean -.041 .074 .026 .071 .127 -.092 .167 .051 .143 .198
Unif -.013 .008 .004 .009 .011 -.014 .020 .010 .017 .019
Zero -.041 .075 .025 .072 .130 -.096 .173 .051 .149 .205
FordB Gauss -.002 .021 -.002 -.026 .030 -.029 .130 .014 -.025 .182
Inv .000 .001 .000 .001 .002 -.007 .017 .007 .011 .021
Opp -.012 .020 .003 .015 .025 -.045 .035 -.022 .026 .034
SubMean -.022 .041 .007 .032 .062 -.071 .090 -.070 .081 .102
Unif -.003 .003 .002 .003 .007 -.009 .008 .004 .003 .007
Zero -.023 .040 .008 .031 .058 -.072 .091 -.069 .081 .103
Wafer Gauss -.060 -.105 .006 .084 -.164 .016 -.098 .119 .177 -.204
Inv .022 .065 .037 .043 .139 -.013 .093 .026 .072 .244
Opp -.020 .012 .013 .011 .021 -.009 .059 .044 .047 .134
SubMean .117 .198 .013 .172 .329 .060 .054 .031 .067 .058
Unif -.011 .012 .035 .019 .034 .030 .024 .066 .023 .102
Zero -.019 .064 .045 .050 .138 .024 .079 .044 .050 .278
ElecDev Gauss -.004 -.010 .009 .006 -.008 .029 -.001 .079 .050 -.036
Inv .014 .059 .080 .052 .071 -.026 .060 .010 .056 .028
Opp .023 .180 .061 .144 .220 .073 .183 .087 .179 .197
SubMean .143 .236 .220 .222 .272 .148 .217 .113 .201 .290
Unif .052 .094 .052 .057 .156 .038 .103 .048 .070 .107
Zero .035 .156 .075 .122 .271 .085 .187 .076 .164 .269
Examining these results in detail, we find that the perturbation strategies exhibit varying levels of performance, with notable dependencies on model architectures and datasets. Zero and SubMean perturbations often achieve the highest DS scores across the different experimental configurations, though Gauss occasionally outperforms them in specific contexts. Importantly, optimal perturbation selection appears highly contingent on model architecture, even when controlling for the dataset. For instance, on the FordB and Wafer datasets, Gauss perturbations perform relatively well for InceptionTime on several attribution methods, but for ResNet, SubMean or Zero yield better results.
We also observe substantial performance variations among the attribution methods. FO, IG and GS tend to achieve the highest DS values across datasets and perturbation strategies, indicating better feature importance identification. While FO frequently demonstrates marginally superior performance compared to IG and GS, this advantage is not consistent across all experimental settings. In contrast, GR shows poor discriminative ability, frequently yielding negative or near-zero DS values, suggesting that its feature importance assignments are often not better than random ordering. Between these extremes, GS typically outperforms GR but falls short of the effectiveness demonstrated by IG, GS and FO, depending on the experimental condition.
@@ 5.2 Distribution Patterns in Attribution Quality
(a) Overall DS distributions per attribution method
(b) Class-specific DS distributions per attribution method
@@
To understand how attribution performance varies between individual instances, we analyze the distributional characteristics of DS scores. For this analysis, we focus on the FordB dataset, InceptionTime architecture, and SubMean perturbation strategy, as these demonstrate patterns typical of our broader findings. Figure 4 presents these distributions, revealing both overall performance patterns and, crucially, class-specific effects.
The aggregate distributions shown in Figure 4(a) reveal notable variation in the effectiveness of attribution methods. IG, GS, and FO exhibit positively skewed distributions with extended tails toward higher DS values, indicating better attribution quality. In contrast, GR and SG demonstrate negative skewness with tails extending toward lower DS values, suggesting less reliable feature identification. Critically, we observe a concentration of scores around zero, implying limited discriminative power in feature ordering for many instances, even among the better-performing methods.
Figure 4(b) presents a class-stratified analysis that uncovers substantial heterogeneity in attribution quality across classes. For IG, GS, and FO, instances from class 1 consistently achieve higher DS scores with pronounced positive tails and minimal negative values. The same methods show markedly different behavior for class 0, where DS scores cluster tightly around zero, indicating minimal perturbation impact. GR and SG show an inverse pattern, with class 1 showing predominantly negative scores while class 0 maintains a more balanced distribution centered near zero.
These pronounced class-dependent variations in attribution performance raise questions about the generalizability of perturbation-based evaluation methods. To determine whether this phenomenon extends beyond our case study, we now examine class-dependent behavior across all experimental conditions.
@@ 5.3 Class-Dependent Effects in Perturbation Analysis
We apply the proposed class-adjusted DS ($\text{DS}_{c}$) metric to all experimental conditions to investigate class-dependent behaviors. Table 4 presents these adjusted scores, which balance average attribution correctness with consistency between classes. Our results show that incorporating class consistency penalties substantially reduces performance metrics across most experimental conditions, particularly for previously high-performing perturbation strategies.
Table 4: Class-adjusted mean degradation scores ($\text{DS}_{c}$ with $\alpha=1$) for named perturbation strategies across datasets, models, and attribution methods. $\text{DS}_{c}$ measures both attribution quality and cross-class consistency of perturbations. Positive values indicate good attribution quality with consistent behavior across classes, while lower values suggest either poor attribution quality or inconsistent performance between classes. The largest values for each dataset and model-attribution column pair are highlighted in bold.
@@
FordA Gauss -.011 .003 .001 -.078 .003 -.008 .001 .000 -.049 .002
Inv -.005 .000 .000 .000 .001 -.007 .000 .000 .000 .000
Opp -.053 .001 .000 .001 .001 -.104 .000 .000 .000 .000
SubMean -.082 .003 .000 .002 .005 -.188 .000 .001 .000 .000
Unif -.025 .000 .000 .000 .000 -.028 .000 .000 .000 .000
Zero -.082 .003 .001 .002 .005 -.197 .000 .001 .000 -.001
FordB Gauss -.004 .003 -.003 -.049 .006 -.068 .015 .009 -.046 .026
Inv -.001 .000 -.001 .000 .000 -.013 .003 .001 .002 .003
Opp -.023 .003 .001 .002 .004 -.091 .004 -.041 .004 .005
SubMean -.044 .006 .000 .005 .010 -.158 .013 -.134 .014 .014
Unif -.007 .001 .000 .001 .001 -.017 .001 .001 .001 .001
Zero -.044 .006 .000 .005 .010 -.159 .013 -.134 .015 .015
Wafer Gauss -.149 -.214 .001 .003 -.367 .003 -.201 -.004 .033 -.419
Inv .008 .003 .000 .002 .002 -.026 .001 -.012 .008 .009
Opp -.042 .002 -.001 .003 .003 -.021 .002 -.011 .010 .008
SubMean .034 .018 -.005 .037 .038 .014 .048 .000 .018 .028
Unif -.020 .001 .001 .002 .001 .002 .002 -.014 .008 .008
Zero -.047 .007 -.001 .011 .013 .012 .009 -.023 .036 .016
ElecDev Gauss -.039 -.036 -.020 -.029 -.044 .011 -.015 .045 .027 -.091
Inv -.024 .012 .047 .009 .025 -.082 .018 -.045 .027 -.036
Opp -.029 .104 .001 .096 .156 .039 .140 .041 .116 .122
SubMean .078 .130 .141 .136 .219 .075 .151 .081 .134 .229
Unif .026 .031 -.003 .020 .082 -.004 .055 .026 .041 .022
Zero -.010 .089 .031 .063 .226 .053 .118 .038 .104 .162
@@
Zero, SubMean and Gauss perturbations, which demonstrated superior performance before, show marked degradation under class-adjusted evaluation, with most DS scores shifting toward zero or even becoming negative. The extent of this impact differs among experimental conditions, particularly across various datasets. Wafer and ElecDev generally show more resilience to class adjustment compared to FordA and FordB, though we still see reduced scores. For ElecDev, the interpretation of these results requires additional context: as the only multiclass dataset in our evaluation, the pairwise averaging of class differences may underestimate class-specific effects due to the higher dimensionality of the classification space.
@@
To examine the mechanisms underlying these class-specific effects, Figure 5 presents a detailed analysis of perturbation impacts across classes on the FordB dataset. Figure 5(a), which focuses on named perturbation strategies with InceptionTime and FO attribution, reveals systematic class-dependent behavior. Class 0 instances exhibit minimal response to perturbation across all strategies, with DS scores tightly clustered around zero, while Class 1 instances demonstrate substantial variability in perturbation response. For Class 1, Gauss is the most effective, followed by Zero and SubMean strategies. This asymmetric response pattern may explain the earlier observed degradation in class-adjusted metrics: perturbation strategies succeed primarily by exploiting class-specific model behaviors.
@@
Analyzing the constant-value perturbations in Figure 5(b) reinforces these findings while providing additional information. The evaluation of perturbation values between -2 and 2 reveals optimal effectiveness at moderate positive values, particularly around 0.5. We now extend the investigation of the constant perturbation strategies to all experimental conditions.
(a) DS for named perturbation strategies
(b) DS for constant perturbation strategies